Help with limiting regent show all steps.


ANAL A.2 Results Weight of crude product obtained: 0.731 Weight of pure product obtained: 0.560 Moles of pure product obtained: 0.0031 mol Melting point of pure product: ) 168-169 (method used - Yield calculation 168.3-1691 i) weight used = 1.003g no. of moles = 0.0073 so LR because leks. Salicylic acid m1 used = 2 weight used = 2.16 no. of moles = 0.021 Acetic anhydride Therefore, the limiting reagent is % Yield = Therefore % yield of pure aspirin = moles of pure product obtained max. moles of product possible x100 % 100 % = x ㅡㅡㅡㅡㅡ ㅡ Appearance of product: Aa
Answer #1

Limiting reagent is salicylic acid because less number of moles.

%yield = 0.0031*100 / 0.0073 = 0.4246*100 %

Therefore % yield of pure aspirin = 42.46 %